OpenStudy (anonymous):

-49x^2 +35x +14 = 0 i tried the formula and ended up with a huge number under the square root. if i complete the square then i will end up with huge fractions. any body have knowledge on this. i will not be permitted to use a calculator on the test. we will only have a pencil and paper.

OpenStudy (anonymous):

do you know the trick how to factorize?

OpenStudy (anonymous):

ax^2+bx+c=0

OpenStudy (anonymous):

i did not focus on factoring because the formula and completing the square yield 100% results for all quadratic problems . i do need to learn how to factor polynomials thought. can you help ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it's very easy

OpenStudy (anonymous):

by which method you need to find it? Completing the square?

OpenStudy (anonymous):

preferably . however, i will use any method that gets the answer quickest. i will only have approximately 1.25 minutes on each problem during the test ...

OpenStudy (anonymous):

completing the square: multiply each side by -1, to get the x^2 positive

OpenStudy (anonymous):

i thought you had to have a coefficient of 1 for A to use this method ..

OpenStudy (anonymous):

so it will be \[49x ^{2}-35x-14\]=0

OpenStudy (anonymous):

(7x-5/2)^2-81/4

OpenStudy (anonymous):

=0

OpenStudy (anonymous):

move 81/4 to the o side, so it will be \[(7x-\frac{ 5 }{ 2 })^{2}=\frac{ 81 }{ 4 }\]

OpenStudy (anonymous):

remove square, so it will be 7x-5/2=9/2

OpenStudy (anonymous):

could you show the way you managed : (7x-5/2)^2-81/4=0 ?

OpenStudy (anonymous):

seems to be you are not enough familiar with completing the square

OpenStudy (anonymous):

i did not focus on factoring when i was learning how to solve quadratics so, i hardly know how to do it.

OpenStudy (anonymous):

do you know the formula of completing the square

OpenStudy (anonymous):

it's not factoring, it is completing the square

OpenStudy (anonymous):

divide b by 2, square the answer and add it to the Ax^2 + Bx the way it was shown to me, was to subtract b from the result and add the difference to both sides of the equation to keep the values down to a manageable size :) so it would work like this is b was 4 and c was 6: 4/2 = 2 2^2 = 4 6-4 = 2 Ax^2 +4x + 6 = 2

OpenStudy (anonymous):

then take a and c to form : (x+6)^2 = 2

OpenStudy (anonymous):

that is what i know about completing the square :)

OpenStudy (anonymous):

oh yeah: x+6=√(2) x=√(2) - 6

OpenStudy (anonymous):

whats that

OpenStudy (anonymous):

the answer, so i had thought.

OpenStudy (anonymous):

you mean of -49x^2 +35x +14 = 0????

OpenStudy (anonymous):

do you want me to explain how completing the square works

OpenStudy (anonymous):

yes, because what i have learned this far is not working :(

OpenStudy (anonymous):

\[(a+b)^{2}=a ^{2}+2ab+b ^{2}\]

OpenStudy (anonymous):

remember?

OpenStudy (anonymous):

yes, i have that in my notes

OpenStudy (anonymous):

so, in many quadratic equations it's not always in that way. so so add and substract additional values

OpenStudy (anonymous):

for example take as (x+2)^2 as our equation

OpenStudy (anonymous):

i took (x+2)^2 to make it simpler

OpenStudy (anonymous):

so \[(x+2)^{2}=x ^{2}+2*2*x+4\]

OpenStudy (anonymous):

and in questions you get problems like x^2+4x+13, and not 4

OpenStudy (anonymous):

i worked out the math once and found that relationship to be true using a calculator, however, i failed to comprehend the relationship.

OpenStudy (anonymous):

so you need to reformulate it in the other form to x^2+4x+4+9

OpenStudy (anonymous):

did you get that?

OpenStudy (anonymous):

4+9=13

OpenStudy (anonymous):

so you can complete the square -----> \[(x+2)^{2}+9\]

OpenStudy (anonymous):

what happened to the 4 ? x^2+4x+13 x^2+4x+4+9 (what happened in between?) (x+2)^2+9

OpenStudy (anonymous):

we know that (x+2)^2=x^2+4x+4

OpenStudy (anonymous):

so the rule of completing the square derived from this

OpenStudy (anonymous):

(x+2)^2 converts to x^2 + 4 . i do not comprehend how the 4x works into this. o_O

OpenStudy (anonymous):

imagine if you were given a question like this \[x^{2}+4x+13\] and asks to complete the square

OpenStudy (anonymous):

you know that \[(x+2)^{2}=x^{2}+4x+4\]

OpenStudy (anonymous):

so you can reformulate it in the form of \[x ^{2}+4x+13=x^{2}+4x+4+9\]

OpenStudy (anonymous):

substitute our square into that

OpenStudy (anonymous):

we will get the new equation of \[(x+2)^{2}+9\]

OpenStudy (anonymous):

so how we can derive the formula using so simple method?

OpenStudy (anonymous):

i guess i will simply have to learn more about factoring polynomials because i am not comprehending this. thanks for trying :)

OpenStudy (anonymous):

in short, formula is like that \[ax ^{2}+bx+c\]

OpenStudy (anonymous):

where \[b=2*\sqrt{a}*\sqrt{c}\]

OpenStudy (anonymous):

\[(\sqrt{a}*x+\frac{ b }{ 2\sqrt{a} })^{2}\]